# A Survey of Vibe Coding with Large Language Models

### v1

@@ 5.1.1 Task Decomposition Strategies
Task decomposition systematically breaks complex problems into manageable subtasks, proven essential from traditional planning to modern LLM applications, enabling coding agents to tackle sophisticated challenges while improving reliability, interpretability, and multi-agent collaboration [70, 348, 72, 73]. Strategies evolved into distinct paradigms including factored cognition for simultaneous processing, process supervision for sequential dependencies, Chain-of-Thought for step-by-step reasoning, and extensions like Tree-of-Thought (ToT) and Graph-of-Thought for multi-path reasoning [72, 74, 349, 75, 201, 76]. Concrete techniques span prompting-based CoT using few-shot or zero-shot instructions [72, 77, 230], Auto-CoT generating samples through clustering [78], Plan-and-Solve devising task-dividing plans [230], tree structures like ToT generating branching explorations [79], and HyperTree Planning enabling divide-and-conquer with Monte Carlo integration [80].
CoT emerges as most popular code generation strategy achieving higher correctness [72, 74], while dynamic strategies like CodePlan employ adaptive algorithms [81, 73], and multi-path approaches simulate generation paths with Monte Carlo optimization, though LLMs potentially hallucinate incoherent plans despite proper instructions [82, 350, 83].
@@ 5.1.2 Plan Formulation Methods
LLMs’ training on natural language and code unlocks reasoning producing structured steps connecting to execution environments, with planning crucial for agents determining action sequences from initial states [73]. Approaches categorize into External Module Augmented Methods integrating planners and memory, Finetuning-based Methods enhancing abilities through trajectory data, and Searching-based Methods identifying optimal solutions [349, 85], while three paradigms emerged: LLM-as-Planner employing inherent reasoning, LLM-as-Facilitator translating to Planning Domain Definition Language (PDDL) [351, 352], and Multi-Agent Planning coordinating multiple agents [73, 252], with task decomposition employing Single-Path following linear trajectories [72], Tree-Based exploring multiple trajectories [79, 353], and Hierarchical Planning at different abstraction levels.
Planning with feedback enables adaptation through ReAct interleaving Thought, Action, and Observation [354, 74], though incremental approaches risk local optima [355], while Memory-Enhanced Planning addresses limitations through Reflexion augmenting agents with Evaluator, Self-reflection, and memory buffers enabling continual improvement [356, 72, 357]. Multi-agent systems coordinate specialized agents distributing planning across components [351, 358], with Integration combining structured PDDL with LLMs where LLMs allocate sub-tasks while planners like Fast Downward generate plans [359, 352, 360, 361]. Challenges persist including Goal Decomposition struggling with actionable steps [362], Long-Horizon Task Management losing coherent strategy [362, 363], and Training Data Scarcity requiring synthetic augmentation [362].
@@ 5.2.1 Overview and Fundamentals
Extending beyond fixed context windows, memory mechanisms emerge as functional phenomena in LLMs demonstrating primacy and recency effects despite lacking explicit subsystems [364, 84, 365]. Contemporary approaches draw parallels to human cognitive architecture, mapping context and parameters to short-term and long-term memory and transforming contextual knowledge into parameter updates [365, 366].
Following this framework, short-term memory constitutes temporary storage with limited capacity, corresponding to context window information accessible during inference [367, 368, 369, 1, 370], while long-term memory stores information over days to years, classified into declarative explicit and non-declarative implicit memory [371]. From an implementation perspective, these memory types classify into parametric memory stored in model weights and non-parametric memory in external systems [371, 372, 373, 85].
@@
CodeAgent [99]
MapCoder [24]
ChatDev [10]
CodeAct [256]
SWE-Search [197]
OpenHands [15]
OpenHands-Versa [23]
MetaGPT [252]
SWE-agent [12]
AutoSafeCoder [374]
AutoCodeRover [13]
Lita [375]
SWE-dev [376]
RepoForge [377]
LessonL [378]
AdaCoder [16]
Code2MCP [379]
ScreenCoder [380]
SimuGen [381]
SoA [194]
SICA [360]
RGD [118]
Guided Code Generation [104]
AgentCoder [104]
AppAgent [382]
Cursor IDE [383]
Claude Code [384]
Gemini Code CLI [385]
Qwen Coder [386]
Codex [387]
@@ 5.2.2 Memory Operations and Management
Accordingly, memory operations enable coding agents to function beyond static context windows, maintaining persistent knowledge about codebases, patterns, and debugging histories [357, 86, 221]. These operations are managed through six fundamental procedures: consolidation, indexing, updating, forgetting, retrieval, and compression [87, 388, 389, 88, 89, 365].
Specifically, the foundation rests on memory reading retrieving relevant information [90], memory writing transforming observations into structured contents, and memory reflection reorganizing knowledge [91, 221, 85]. Comprehensive systems implement specialized functions including consolidation, indexing, and updating employing locate-and-edit techniques [390, 92, 14, 389, 391, 89], with optimization extracting meta-insights and selective management yielding 10% gains [388, 392].
@@ 5.2.3 Memory Architecture in Coding Agents
In practice, short-term memory plays critical role in hierarchical architectures, enabling immediate feedback integration demonstrated by FALCON’s two-level structure [393, 357], while dynamic management with MemTool achieves 90-94% efficiency [394].
In contrast, long-term memory implementations utilize external memory-based methods with physical storage maintaining historical information. Vector databases emerged as prominent non-parametric solution achieving long-term effects and allowing domain-specific knowledge access [395, 396, 397], though struggling with belief updating challenges. To address these limitations, Memory Bank stores memory with temporal timestamps and implements Ebbinghaus-inspired forgetting curves [89, 398, 396, 85]. These implementations categorize into three architectural patterns: Long-Context Agents with First-In-First-Out (FIFO) eviction, RAG Agents using external pools, and Agentic Memory Agents employing iterative reasoning [399, 365]. Furthermore, contemporary architectures adopt dual-component systems mirroring human cognition through frameworks like RAISE [400] and InfiniteICL [401], transforming temporary knowledge into permanent updates. Additionally, specialized mechanisms demonstrate distinct optimization strategies. Think-in-Memory eliminates redundant reasoning by caching intermediate results [402], whereas MemGPT leverages virtual context management to extend effective context windows through intelligent memory paging [357], thus addressing belief updating challenges inherent in static vector database approaches.
@@ 5.3.1 Tool Invocation
Tool integration with LLMs fundamentally transformed AI-powered development [93, 74], with Tool-Augmented Models extending functionalities by integrating external tools [94, 95, 74], producing AI agentic programming where LLM-based agents autonomously plan, execute, and interact with compilers, debuggers, and version control reshaping development practices. Early task-specific systems suffered limited scalability [97], while ReAct established observation-action-thought pattern becoming foundation for LangChain and AutoGen [97, 74], with modern frameworks adopting standardized paradigms decomposing capabilities into specialized components [403] and enabling automatic transformation of documentation into callable functions [96, 98].
Tool invocation evolved beyond traditional JSON-based approaches [74, 404], with executable code-based frameworks [256, 127] consolidating actions into unified Python space enabling dynamic revision and composition, demonstrating substantial improvements over predefined constraints. Systems like ToolCoder reduced hallucination errors [243, 93, 252], while agents integrating specialized tools [99] demonstrated significant improvements outperforming commercial products like GitHub Copilot [243]. Protocol standardization through Model Context Protocol established schema-driven interfaces, with advanced discovery mechanisms [97, 101] enabling dynamic selection through iterative invocation and document-weighted averaging addressing scalability across large-scale repositories. Enterprise-focused frameworks [405, 406, 102, 73] automated pipelines transforming API specifications into executable Python functions, substantially enhancing accuracy while eliminating scaffolding efforts, with reinforcement learning integration [407, 116] unifying tool usage with reasoning through policy optimization.
@@
Embedded Cursor [383] AI-native code editor with multi-file editing, intelligent refactoring, and conversational AI VS Code fork with deep AI integration, composer agent mode, supports Claude/GPT-4/Gemini models, codebase-wide context awareness Complex projects requiring intelligent refactoring, full-stack development, professional developers
@@
Copilot [408] Real-time code completion, generation, and explanation across multiple IDEs Trained on GitHub repositories, supports VS Code/JetBrains/Neovim, inline suggestions, chat interface Daily coding assistance, autocomplete-focused workflows, GitHub-integrated development
@@
Windsurf [409] Agentic IDE with autonomous coding capabilities and deep contextual understanding Cascade Flow system, auto-executes terminal commands, supports multiple AI models, built on VS Code Large codebases, multi-file reasoning, team collaboration with advanced workflow automation
@@
Tabnine [410] Enterprise-grade AI code completion with privacy-first approach Local/on-premise deployment options, trained only on permissive licenses, custom model training on proprietary code Enterprise environments, regulated industries (finance, healthcare), privacy-sensitive projects
@@
CodeWhisperer [411] AWS-optimized code generation and security scanning Integrated with AWS services, reference tracking for license compliance, security vulnerability scanning AWS cloud development, serverless applications, cloud-native architecture
@@
End-to-End v0 [412] Natural language to React/Next.js UI generation with deployment Generates production-ready React + Tailwind code, integrated security checks, one-click Vercel deployment Frontend prototyping, UI/UX design-to-code, rapid MVP development
@@
Replit [413] Cloud-based autonomous app builder from natural language descriptions Complete environment setup, integrated databases and authentication, multi-language support, browser-based testing Non-technical users, rapid prototyping, educational projects, small business tools
@@
Bolt.new [414] Browser-native full-stack development with instant preview and deployment WebContainers technology, runs Node.js in browser, supports npm packages, StackBlitz infrastructure Full-stack prototyping, quick demos, cloud-native development without local setup
@@
CLI Claude Code [384] Terminal-based agentic coding tool with autonomous capabilities Powered by Claude Sonnet 4.5, local-first execution, MCP integration, checkpoint/rewind system, VS Code extension Complex terminal workflows, autonomous code refactoring, multi-file operations
@@
Gemini CLI [415] Open-source AI agent for terminal-native development Gemini 2.5 Pro with 1M token context, agent mode with multi-step planning, free tier (60 RPM, 1000 requests/day) Terminal-first developers, Google Cloud workflows, budget-conscious teams
@@
Qwen Code [416] Research-focused CLI tool for agentic coding tasks Powered by Qwen3-Coder models (480B MoE), 256K-1M token context, enhanced parser Open-source enthusiasts, researchers, privacy-focused local deployments
@@
Aider [417] Open-source AI pair programming in terminal Multi-LLM support (Claude, GPT-4o, DeepSeek), codebase mapping, automatic Git commits, voice commands Budget-conscious developers, multi-model workflows, Git-centric development
@@
Codex [387] Cloud-based software engineering agent that writes features, fixes bugs, answers codebase questions, and proposes PRs in isolated sandboxes Parallel task execution; per-task cloud sandboxes; CLI & IDE extension; GitHub PR (@codex) integration; configurable via ˜/.codex/config; MCP tool access Pro teams automating code reviews, refactors, tests, and on-call maintenance across large repositories
@@
Cline [418] Open-source coding agent in IDE/CLI that understands whole codebases, plans complex changes, and executes multi-step tasks with user approval Client-side, zero-trust architecture; model-agnostic (Claude/Gemini/etc.); reads/writes files, runs terminal commands & browser; MCP integration; SOC 2/GDPR; enterprise team/billing options Enterprise/security-sensitive orgs; large-codebase refactors; multi-model workflows needing transparent, auditable control

### v2

@@ 5.1.1 Task Decomposition Strategies
Task decomposition systematically breaks complex problems into manageable subtasks, proven essential from traditional planning to modern LLM applications, enabling coding agents to tackle sophisticated challenges while improving reliability, interpretability, and multi-agent collaboration [stiennon2020learning, huang2024levels, wei2022chain, hu2024agentgen]. Strategies evolved into distinct paradigms including factored cognition for simultaneous processing, process supervision for sequential dependencies, Chain-of-Thought for step-by-step reasoning, and extensions like Tree-of-Thought (ToT) and Graph-of-Thought for multi-path reasoning [wei2022chain, yao2022react, cao2025large, lu2025decompose, shinn2023reflexion, shen2023hugginggpt]. Concrete techniques span prompting-based CoT using few-shot or zero-shot instructions [wei2022chain, kojima2022large, zhao2023depth], Auto-CoT generating samples through clustering [zhang2022automatic], Plan-and-Solve devising task-dividing plans [zhao2023depth], tree structures like ToT generating branching explorations [yao2023tree], and HyperTree Planning enabling divide-and-conquer with Monte Carlo integration [gui2025hypertree].
CoT emerges as most popular code generation strategy achieving higher correctness [wei2022chain, yao2022react], while dynamic strategies like CodePlan employ adaptive algorithms [wen2024codeplan, hu2024agentgen], and multi-path approaches simulate generation paths with Monte Carlo optimization, though LLMs potentially hallucinate incoherent plans despite proper instructions [tian2024sqlucid, jayagopal2022exploring, patel2022question].
@@ 5.1.2 Plan Formulation Methods
LLMs’ training on natural language and code unlocks reasoning producing structured steps connecting to execution environments, with planning crucial for agents determining action sequences from initial states [hu2024agentgen]. Approaches categorize into External Module Augmented Methods integrating planners and memory, Finetuning-based Methods enhancing abilities through trajectory data, and Searching-based Methods identifying optimal solutions [cao2025large, zhang2025multiple], while three paradigms emerged: LLM-as-Planner employing inherent reasoning, LLM-as-Facilitator translating to Planning Domain Definition Language (PDDL) [molinari2025towards, helmert2006fast], and Multi-Agent Planning coordinating multiple agents [hu2024agentgen, hong2024metagpt], with task decomposition employing Single-Path following linear trajectories [wei2022chain], Tree-Based exploring multiple trajectories [yao2023tree, besta2023graph], and Hierarchical Planning at different abstraction levels.
Planning with feedback enables adaptation through ReAct interleaving Thought, Action, and Observation [pitkranta2025hada, yao2022react], though incremental approaches risk local optima [chen2025enhancing], while Memory-Enhanced Planning addresses limitations through Reflexion augmenting agents with Evaluator, Self-reflection, and memory buffers enabling continual improvement [luo2025large, wei2022chain, packer2023memgpt]. Multi-agent systems coordinate specialized agents distributing planning across components [molinari2025towards, ouyang2024long], with Integration combining structured PDDL with LLMs where LLMs allocate sub-tasks while planners like Fast Downward generate plans [zhang2024lamma, helmert2006fast, singh2024twostep, sun2023pearl]. Challenges persist including Goal Decomposition struggling with actionable steps [erdogan2025plan], Long-Horizon Task Management losing coherent strategy [erdogan2025plan, kienle2025learning], and Training Data Scarcity requiring synthetic augmentation [erdogan2025plan].
@@ 5.2.1 Overview and Fundamentals
Extending beyond fixed context windows, memory mechanisms emerge as functional phenomena in LLMs demonstrating primacy and recency effects despite lacking explicit subsystems [janik2023aspects, vaswani2017attention, shan2025cognitive]. Contemporary approaches draw parallels to human cognitive architecture, mapping context and parameters to short-term and long-term memory and transforming contextual knowledge into parameter updates [shan2025cognitive, parthasarathy2024ultimate].
Following this framework, short-term memory constitutes temporary storage with limited capacity, corresponding to context window information accessible during inference [ren2024brain, lu2023towards, xie2024travelplanner, brown2020language, zhang2023igniting], while long-term memory stores information over days to years, classified into declarative explicit and non-declarative implicit memory [he2024human]. From an implementation perspective, these memory types classify into parametric memory stored in model weights and non-parametric memory in external systems [he2024human, wu2025from, han2024llm, zhang2025multiple].
@@
CodeAgent [zhang2024codeagent]
MapCoder [zhao2025mage]
ChatDev [qian2023chatdev]
CodeAct [wang2024executable]
SWE-Search [antoniades2024swe]
OpenHands [wang2024openhands]
OpenHands-Versa [ho2025verilogcoder]
MetaGPT [hong2024metagpt]
SWE-agent [yang2024swe]
AutoSafeCoder [witt2025open]
AutoCodeRover [zhang2024autocoderover]
Lita [madison2024scalable]
SWE-dev [wang2025swe]
RepoForge [chen2025repoforge]
LessonL [liu2025lessons]
AdaCoder [robeyns2025self]
Code2MCP [ouyang2025transforming]
ScreenCoder [jiang2025screencoder]
SimuGen [ren2025simugen]
SoA [ishibashi2024self]
SICA [singh2024twostep]
RGD [jin2024rgd]
Guided Code Generation [huang2023agentcoder]
AgentCoder [huang2023agentcoder]
AppAgent [zhang2025appagent]
Cursor IDE [cursor-ide]
Claude Code [claude-code]
Gemini Code CLI [gemini-code-cli]
Qwen Coder [qwen-coder]
Codex [codex]
@@ 5.2.2 Memory Operations and Management
Accordingly, memory operations enable coding agents to function beyond static context windows, maintaining persistent knowledge about codebases, patterns, and debugging histories [packer2023memgpt, xu2025mem, zhang2025survey]. These operations are managed through six fundamental procedures: consolidation, indexing, updating, forgetting, retrieval, and compression [zhang2025memengine, xiong2025memory, du2025rethinking, feng2024tasl, zhong2023memorybank, shan2025cognitive].
Specifically, the foundation rests on memory reading retrieving relevant information [park2023generative], memory writing transforming observations into structured contents, and memory reflection reorganizing knowledge [huang2022language, zhang2025survey, zhang2025multiple]. Comprehensive systems implement specialized functions including consolidation, indexing, and updating employing locate-and-edit techniques [hou2024my, modarressi2023ret, xia2024agentless, du2025rethinking, bae2022keep, zhong2023memorybank], with optimization extracting meta-insights and selective management yielding 10% gains [xiong2025memory, chhikara2025mem0].
@@ 5.2.3 Memory Architecture in Coding Agents
In practice, short-term memory plays critical role in hierarchical architectures, enabling immediate feedback integration demonstrated by FALCON’s two-level structure [zhan2025coderagent, packer2023memgpt], while dynamic management with MemTool achieves 90-94% efficiency [lumer2025memtool].
In contrast, long-term memory implementations utilize external memory-based methods with physical storage maintaining historical information. Vector databases emerged as prominent non-parametric solution achieving long-term effects and allowing domain-specific knowledge access [lewis2020retrieval, kynoch2023recallm, wu2022memorizing], though struggling with belief updating challenges. To address these limitations, Memory Bank stores memory with temporal timestamps and implements Ebbinghaus-inspired forgetting curves [zhong2023memorybank, xing2025structured, kynoch2023recallm, zhang2025multiple]. These implementations categorize into three architectural patterns: Long-Context Agents with First-In-First-Out (FIFO) eviction, RAG Agents using external pools, and Agentic Memory Agents employing iterative reasoning [hu2025evaluating, shan2025cognitive]. Furthermore, contemporary architectures adopt dual-component systems mirroring human cognition through frameworks like RAISE [liu2024llm] and InfiniteICL [cao2025infiniteicl], transforming temporary knowledge into permanent updates. Additionally, specialized mechanisms demonstrate distinct optimization strategies. Think-in-Memory eliminates redundant reasoning by caching intermediate results [pawar2024what], whereas MemGPT leverages virtual context management to extend effective context windows through intelligent memory paging [packer2023memgpt], thus addressing belief updating challenges inherent in static vector database approaches.
@@ 5.3.1 Tool Invocation
Tool integration with LLMs fundamentally transformed AI-powered development [schick2023toolformer, yao2022react], with Tool-Augmented Models extending functionalities by integrating external tools [chen2023large, shen2024small, yao2022react], producing AI agentic programming where LLM-based agents autonomously plan, execute, and interact with compilers, debuggers, and version control reshaping development practices. Early task-specific systems suffered limited scalability [fei2025mcp], while ReAct established observation-action-thought pattern becoming foundation for LangChain and AutoGen [fei2025mcp, yao2022react], with modern frameworks adopting standardized paradigms decomposing capabilities into specialized components [shen2024llm] and enabling automatic transformation of documentation into callable functions [hou2025model, shi2024tool].
Tool invocation evolved beyond traditional JSON-based approaches [yao2022react, qin2023tool], with executable code-based frameworks [wang2024executable, islam2024mapcoder] consolidating actions into unified Python space enabling dynamic revision and composition, demonstrating substantial improvements over predefined constraints. Systems like ToolCoder reduced hallucination errors [dong2025survey, schick2023toolformer, hong2024metagpt], while agents integrating specialized tools [zhang2024codeagent] demonstrated significant improvements outperforming commercial products like GitHub Copilot [dong2025survey]. Protocol standardization through Model Context Protocol established schema-driven interfaces, with advanced discovery mechanisms [fei2025mcp, lumer2025scalemcp] enabling dynamic selection through iterative invocation and document-weighted averaging addressing scalability across large-scale repositories. Enterprise-focused frameworks [agarwal2025automated, mastouri2025making, ni2025doc2agent, hu2024agentgen] automated pipelines transforming API specifications into executable Python functions, substantially enhancing accuracy while eliminating scaffolding efforts, with reinforcement learning integration [singh2025agentic, jin2025reveal] unifying tool usage with reasoning through policy optimization.
@@
Embedded Cursor [cursor-ide] AI-native code editor with multi-file editing, intelligent refactoring, and conversational AI VS Code fork with deep AI integration, composer agent mode, supports Claude/GPT-4/Gemini models, codebase-wide context awareness Complex projects requiring intelligent refactoring, full-stack development, professional developers
@@
Copilot [copilot] Real-time code completion, generation, and explanation across multiple IDEs Trained on GitHub repositories, supports VS Code/JetBrains/Neovim, inline suggestions, chat interface Daily coding assistance, autocomplete-focused workflows, GitHub-integrated development
@@
Windsurf [windsurf] Agentic IDE with autonomous coding capabilities and deep contextual understanding Cascade Flow system, auto-executes terminal commands, supports multiple AI models, built on VS Code Large codebases, multi-file reasoning, team collaboration with advanced workflow automation
@@
Tabnine [tabnine] Enterprise-grade AI code completion with privacy-first approach Local/on-premise deployment options, trained only on permissive licenses, custom model training on proprietary code Enterprise environments, regulated industries (finance, healthcare), privacy-sensitive projects
@@
CodeWhisperer [CodeWhisperer] AWS-optimized code generation and security scanning Integrated with AWS services, reference tracking for license compliance, security vulnerability scanning AWS cloud development, serverless applications, cloud-native architecture
@@
End-to-End v0 [v0] Natural language to React/Next.js UI generation with deployment Generates production-ready React + Tailwind code, integrated security checks, one-click Vercel deployment Frontend prototyping, UI/UX design-to-code, rapid MVP development
@@
Replit [replit] Cloud-based autonomous app builder from natural language descriptions Complete environment setup, integrated databases and authentication, multi-language support, browser-based testing Non-technical users, rapid prototyping, educational projects, small business tools
@@
Bolt.new [bolt] Browser-native full-stack development with instant preview and deployment WebContainers technology, runs Node.js in browser, supports npm packages, StackBlitz infrastructure Full-stack prototyping, quick demos, cloud-native development without local setup
@@
CLI Claude Code [claude-code] Terminal-based agentic coding tool with autonomous capabilities Powered by Claude Sonnet 4.5, local-first execution, MCP integration, checkpoint/rewind system, VS Code extension Complex terminal workflows, autonomous code refactoring, multi-file operations
@@
Gemini CLI [googlecli] Open-source AI agent for terminal-native development Gemini 2.5 Pro with 1M token context, agent mode with multi-step planning, free tier (60 RPM, 1000 requests/day) Terminal-first developers, Google Cloud workflows, budget-conscious teams
@@
Qwen Code [qwencode] Research-focused CLI tool for agentic coding tasks Powered by Qwen3-Coder models (480B MoE), 256K-1M token context, enhanced parser Open-source enthusiasts, researchers, privacy-focused local deployments
@@
Aider [aider] Open-source AI pair programming in terminal Multi-LLM support (Claude, GPT-4o, DeepSeek), codebase mapping, automatic Git commits, voice commands Budget-conscious developers, multi-model workflows, Git-centric development
@@
Codex [codex] Cloud-based software engineering agent that writes features, fixes bugs, answers codebase questions, and proposes PRs in isolated sandboxes Parallel task execution; per-task cloud sandboxes; CLI & IDE extension; GitHub PR (@codex) integration; configurable via ˜/.codex/config; MCP tool access Pro teams automating code reviews, refactors, tests, and on-call maintenance across large repositories
@@
Cline [cline] Open-source coding agent in IDE/CLI that understands whole codebases, plans complex changes, and executes multi-step tasks with user approval Client-side, zero-trust architecture; model-agnostic (Claude/Gemini/etc.); reads/writes files, runs terminal commands & browser; MCP integration; SOC 2/GDPR; enterprise team/billing options Enterprise/security-sensitive orgs; large-codebase refactors; multi-model workflows needing transparent, auditable control